# FlipedRAG: Black-Box Opinion Manipulation Attacks to Retrieval-Augmented Generation of Large Language Models

### v1

@@ 4.2 Experiment Details
Input: target black-box RAG model $\textit{RAG}_{\text{black}}$, target retrieval model $RM$, surrogate model $\boldsymbol{M_{i}}$, controversial topics $\boldsymbol{Q}$, target topic $\boldsymbol{q}$, target opinion $S_{t}$, corpus $Docs$ , target documents with target opinion $Docs_{t}$, target document $\boldsymbol{d_{t}}$, relevant document $d_{+}$, random sampled document $d_{-}$
@@
$i_{1}$ = "—Here is the user question—
@@
Please COPY all the given context altogether in [[ ]] including all marks and symbols. Do not omit any sentence of the context."
$i_{2}$ = "Use the following pieces of retrieved context to answer the question…"
@@
$\boldsymbol{OpinionClassify}$: Classify the opinion of the content into "support", "neutral" or "oppose".
@@ 3       INIT: Dataset $\mathcal{D}\leftarrow\{\}$
             // $\textit{RAG}_{\text{black}}(\boldsymbol{q}_{m}\oplus i_{1})$approx $RM(\boldsymbol{q}_{m})$
@@ 15       INIT: RAG Response Set $\boldsymbol{Res}\leftarrow\{\}$
17             rank list $\boldsymbol{R}_{m}\leftarrow$ $\boldsymbol{M_{i}(\boldsymbol{q}_{m};Docs)}$
18             $\boldsymbol{anchor}_{m}\leftarrow$ top-1($\boldsymbol{R}_{m}$)
19             for $d_{j}\in\boldsymbol{R}_{m}$ do
20                   if $\boldsymbol{OpinionClassify}(d_{j})=S_{t}$ then
21                         $\boldsymbol{Docs_{t}}\leftarrow$ $d_{j}$
22
@@ 27            $\boldsymbol{Res}\leftarrow$ $\textit{RAG}_{\text{black}}(\boldsymbol{q}_{m};\boldsymbol{Docs})$
Algorithm 1 FlipedRAG (black-box)
4.3 Research Questions
We propose three research questions to evaluate the effect of our method about black-box retriever imitation and opinion manipulation in RAG.
RQ2: How significant does opinion manipulation influence the response of the LLM in RAG?
4.4 Evaluation Metrics
For RQ1, we aim to evaluate the relevance ranking ability of the surrogate model trained through black-box imitation of $\text{RAG}_{\text{black}}$ on the standard ranking task and assess the similarity between their ranking performance and that of the target retrieval model. Additionally, when adversarial triggers are applied to the target retrieval model, we compare the manipulation effect of triggers generated from the surrogate model versus those generated from the original BERT model, which is not trained by black-box imitation. If the trigger generated based on the surrogate model shows better manipulation performance on the ranking of the target retrieval model, it indicates that the surrogate model has learned internal knowledge features of the target retrieval model through black-box retriever imitation training.
So we use Mean Reciprocal Rank (MRR) and Normalized Discounted Cumulative Gain (NDCG) to reflect the ranking ability of the models; higher values of these metrics indicate stronger ranking ability of the model in terms of relevance. Then we use Inter Ranking Similarity (Inter) and Rank Biased Overlap (RBO) to measure the similarity between the ranking performance of the surrogate model and that of the target retrieval model.
Mean Reciprocal Rank (MRR). MRR reflects the effectiveness of ranking by calculating the reciprocal of the rank at which the first relevant item appears. So MRR highlights relevant items. The MRR value is averaged over multiple queries and it ranges from 0 to 1, higher MRR values indicate better rankings with relevant documents found earlier in the ranking. MRR@10 means the score only takes the first 10 items in the ranking into account, the same below.
Normalized Discounted Cumulative Gain (NDCG). NDCG measure first calculate Cumulative Gain (CG), the sum of relevance scores for the results in the ranking. Later, considering that relevant items appearing earlier in the ranking are more valuable, it adjust CG to Discounted Cumulative Gain (DCG) by accounting for the position of each candidate item. At last, it normalizes the actual DCG with the maximum possible DCG that could be achieved if all relevant items are ranked in perfect order, resulting the NDCG score ranging from 0 to 1. An NDCG score close to 1 indicates that the ranking is nearly ideal, with highly relevant items near the top of the list.
Rank Biased Overlap (RBO). RBO considers both the positions of candidate items and the extent of overlap between the two rankings and it ranges from 0 to 1, too. RBO@1000 calculates the top 1000 weighted ranking overlap.
For RQ2, we aim to evaluate the manipulation performance of our attack on retriever rankings and LLM responses. So we firstly evaluate the manipulation performance with the average proportion of target opinions in the top 3 rankings before and after manipulation (Top3origin, Top3attacked). Based on Top3origin and Top3attacked, we calculate the difference Top3var between these two values to reflect the extent of change in the ranking manipulation of candidate documents holding the target opinion $S_{t}$. Additional, we also use Ranking Attack Success Rate (RASR), Boost Rank (BRank) and Boost Rate (BRate) to reflect how effective our attack distort the retrieval ranking of RAG. Here are the details of the metrics on ranking manipulation above.
@@ 4.4 Evaluation Metrics
Top3v. Top3v is obtained by subtracting Top3origin from Top3attacked. Note that Top3v ranges from -1 to 1, and a positive value indicates that the manipulation successfully increased the proportion of the target opinion in the top-3 ranking, while a negative value suggests a detrimental effect of the manipulation. The higher the Top3v value, the more effective the manipulation. Top6v is also defined as above.
@@
Ranking Attack Success Rate (RASR). RASR measures the average rate of successfully boosted candidates in the ranking of each query. The higher is the RASR value, the more successful is the ranking attack.
@@
Boost Rank (BRank):BRank is the average of the total rank improvements for all target documents under each query. Note that a target document $d_{t}$ contributes a negative value to BRank if its rank is reduced accidentally by the attack.
Apart from the evaluation on ranking manipulation, we use Opinion Manipulation Success Rate (OMSR) and Average Stance Variation (ASV) to evaluate the significance of our opinion manipulation on the response of the LLM in black-box RAG.
@@
Opinion Manipulation Success Rate (OMSR). OMSR measures the average rate of successfully manipulated LLM responses at the opinion level. We consider one manipulation as a success if it guides the LLM to produce a response whose opinion is closer to the target opinion $S_{t}$ compared with the original response without RAG attack. Given the target opinion "support"(Pro), successful opinion manipulation can be from "oppose"(Con) to "neutral", from "oppose" to "support", and from "neutral" to "support". Given the target opinion "oppose", successful opinion manipulation can be from "support" to "neutral", from "support" to "oppose", and from "neutral" to "oppose".
@@
Average Stance Variation (ASV). ASV represents the average increase of opinion scores of LLM responses in the direction of the target opinion $S_{t}$ after manipulation. In this research, we assign a value of 0 to opposing stance, 1 to neutral stance, and 2 to supporting stance. A positive ASV indicates that the opinion manipulation towards $S_{t}$ is effective, while a negative ASV indicates that the manipulation actually makes the opinions of RAG responses deviate from $S_{t}$. ASV reflects how much the shift in opinion polarity is. The larger the ASV value, the more successful the opinion manipulation of RAG responses.
5 Experimental Results Analysis
5.1 RQ1: Does black-box retriever imitation effectively learn about the internal knowledge of the retriever of RAG?
After imitating the retrieval model of $\text{RAG}_{\text{black}}$ to obtain the surrogate model, we first compare the ranking ability of the surrogate model $M_{i}$ and the target retrieval model $RM$, as well as the similarity of their ranking results to ensure that the surrogate model has learned the capabilities of the black-box retrieval model.
TABLE I: Comparison(%) of ranking ability between the surrogate model and the target retrieval model(based on the target retrieval model).
Model MRR@10 NDCG@10 Inter@10 RBO@10
Contriever 79.06 67.71
$M_{Contriever}$ 86.82 69.96 62.55 48.52
Co-Condenser 87.01 68.16
$M_{CoCondenser}$ 84.06 67.98 61.86 49.40
ANCE 85.34 67.84
$M_{ANCE}$ 83.79 66.49 56.74 45.66
As shown in Table 1, the surrogate model $M_{i}$ trained by black-box imitation is similar to the target retrieval model in relevance ranking ability for their close NDCG@10 scores. Their ranking results are also similar enough, validating the effectiveness of the black-box imitation. For convenience, we use $M_{Contriever}$ to represent the surrogate model that has its training data collected from $\text{RAG}_{\text{black}}$ which uses Contriever as the retrieval model. The method of representation applies to surrogate models based on ranking data of Co-Condenser and ANCE as well.
In addition to relevance ranking performance, we place great importance on the manipulation performance of the adversarial triggers generated from the surrogate model $M_{i}$, because this performance also reflects the effectiveness of black-box imitation. Assuming we produce adversarial triggers by the white-box target retrieval model, its manipulation effect on the target retrieval model should be significantly better than the that of triggers generated by the BERT model for transfer attacks. The ranking manipulation effect of the surrogate model $M_{i}$ on the target retrieval model should fall between the manipulation effect of the white-box target retrieval model and that of BERT-based transfer attacks as $M_{i}$ has acquired some of the target model’s capabilities. In other words, only if triggers from the surrogate model has the better ranking manipulation effect on the target retrieval model than that from BERT, we consider the black-box imitation process to be successful. So we take BERT as the baseline model to evaluate the manipulation effect of black-box imitation process.
TABLE II: Comparison of Manipulation Effect between the Surrogate Model $M_{i}$ and the Baseline Model.
$Top3_{v}$ $Top6_{v}$ $BRank$ $RASR\%$ $OMSR\%$ $ASV$
Contriever PRO baseline 0.31 0.28 72.33 70.73 43.33 0.50
surrogate 0.37 0.38 111.03 74.22 63.33 0.70
CON baseline 0.27 0.24 67.1 71.93 40.00 0.27
surrogate 0.37 0.35 105.13 78.26 50.00 0.47
Co-Condenser PRO baseline 0.22 0.12 43.53 58.91 23.33 0.20
surrogate 0.27 0.22 63.70 62.01 43.33 0.27
CON baseline 0.07 0.13 37.43 57.11 20.00 0.07
surrogate 0.17 0.19 58.96 64.42 33.33 0.27
ANCE PRO baseline 0.30 0.22 57.20 62.79 43.33 0.43
surrogate 0.38 0.32 85.96 69.76 53.33 0.47
CON baseline 0.24 0.19 50.90 64.42 30.00 0.20
surrogate 0.33 0.32 89.40 74.50 36.66 0.37
As shown in Table 2, we find that in terms of ranking manipulation effectiveness, the surrogate models after black-box imitation outperform the baseline models. In both ranking attack success rate (RASR) and boosted rank (BRank) of target documents, triggers generated from the surrogate model achieve better transfer attack effects on the target retrieval model. Specifically, $M_{Contriever}$ has the best ranking manipulation performance on the triggers generated for transfer attacking, especially in terms of boost rank and $Top3_{v}$. $M_{ANCE}$ demonstrates ranking manipulation effectiveness comparable to that of $M_{Contriever}$, whereas the manipulation effect achieved by $M_{CoCondenser}$ is less ideal by comparison. It indicates that Co-Condenser, as a dense retrieval model, is more robust and reliable than Contriver and ANCE.
Additionally, the surrogate models, compared to the baseline BERT, promote more documents holding the target opinion into top 3 range, thereby having a significant impact on the response of $\text{RAG}_{\text{black}}$. Thus, we can see that the transfer attack based on the surrogate model achieves stronger opinion manipulation by successfully flipping opinions of RAG response in the direction of target opinion. Black-box imitation enables the surrogate model to learn the features of the black-box retrieval model, allowing triggers generated by the surrogate model to induce a greater shift in opinion polarity in LLMs’ response compared to the baseline model. Table 2 also indicates that adversarial triggers generated by $M_{Contriever}$ and $M_{ANCE}$ place more target documents in top 3 range. with greater extent of rank promoting than the baseline model has. As a result, $M_{Contriever}$ and $M_{ANCE}$ achieve relatively better performance on OMSR and ASV.
It is interesting that these dense retrieval models seem to favor supportive opinions. They exhibit better ranking manipulation capability on supportive target opinion than opposing target opinion.
In short, the above experiments demonstrate that the surrogate model obtained by black-box imitation not only achieves relevance ranking performance similar to that of the target retrieval model but also generates adversarial triggers with better ranking manipulation effects compared to the baseline. This indicates that the black-box imitation process is able to help BERT learn internal knowledge about the target black-box retrieval model.
5.2 RQ2: How significant does opinion manipulation influence the response of the LLM in RAG?
After the black-box imitation training, we conduct opinion manipulation experiments across different $\text{RAG}_{\text{black}}$ (based on Llama-3, Mixtral, or Vicuna-v1.5) with adversarial triggers generated by the three white-box surrogate models.
TABLE III: RAG Opinion Manipulation Effect among Different LLMs.
Llama-3 Mixtral Vicuna Llama-3 Mixtral Vicuna
$M_{Contriever}$ PRO 63.33 46.66 36.66 0.70 0.57 0.50
CON 50.00 46.66 46.66 0.47 0.40 0.60
$M_{CoCondenser}$ PRO 43.33 36.66 36.66 0.30 0.37 0.43
CON 33.33 36.66 33.33 0.27 0.37 0.33
$M_{ANCE}$ PRO 53.33 56.66 46.66 0.47 0.53 0.53
CON 36.66 40.00 46.66 0.37 0.47 0.60
As shown in Table 3, our transfer attack with black-box imitation achieve significant opinion manipulation effect on all three LLMs. As a whole, it has a 40% – 50% chance of successful opinion polarity flipping and realizes about 0.46 in opinion polarity shift of range 0 to 2. Moreover, table 3 reveals that FlipedRAG attacks based on the surrogate models $M_{Contriever}$ and $M_{ANCE}$ are more effective in corrupting the RAG corpus and manipulating the RAG opinion with relatively higher OMSR and ASV value. That is because these attacks exhibit superior performance on ranking distortion, especially in $Top3_{v}$ aspect.
From the perspective of opinion polarity manipulation, we observe that the impact of the ranking distortion caused by transfer attacks varies in opinion manipulation effectiveness across different LLMs. Llama-3 and Mixtral have higher OMSR score in contrast to Vicuna for their greater instruction-following ability. Thus, Llama-3 and Mixtral are more susceptible to the adversarial documents promoted to top 3 of the ranking, generating responses based on the stance presented in the context that is composed of these adversarial documents. On the contrary, for ASV, Vicuna actually demonstrates a relatively greater shift in opinion polarity. We analyze that this is due to the better diversity in responses from Llama-3 and Mixtral when answering questions, resulting in less stable opinion polarity in the generated responses. In some cases, opinion manipulation is unsuccessful, instead of turning the opinion of $\text{RAG}_{\text{black}}$ response to the target opinion or make it remain unchanged, manipulation causes the opinion polarity to shift in the opposite direction of the target opinion, thereby lowering the ASV value. The cause of these reverse manipulation results may also stem from the randomness of LLMs generation, although this influence is relatively minor.
To further investigate the fluctuations in opinions resulting from the generation randomness inherent in LLMs, we query the the RAG system twice, which uses Contriver as its retrieval model and has a clean database without manipulation, to get two responses on the same topic. The results are illustrated in the Table 4. We can observe that the responses generated by LLMs on controversial topics do exhibit fluctuations in opinions. However, these fluctuations are minimal so that the change of the text opinion polarity is significantly less than that resulting from deliberate manipulation attacks.
TABLE IV: Opinion Fluctuation of RAG Response on Clean Database.
LLM Target Opinion OMSR(%) ASV
Llama-3 PRO 20 0.17
CON 10 0.0
Mixtral PRO 10 0.07
CON 20 0.10
Vicuna PRO 13.3 -0.03
CON 16.7 0.07
Figure 5: Overall effect of RAG opinion manipulation on three LLMs under black-box setting (retrieval model: Contriever).
In addition, we observe that LLMs are more susceptible to manipulation in the direction of supportive target opinion. This is likely due to the dense retrieval model, acting as the retriever, favoring supportive stances.
From the perspective of overall opinion manipulation of $\text{RAG}_{\text{black}}$ response, as shown in Figure 5, our transfer attack with black-box imitation is able to significantly alter the opinion of RAG-generated content in a specified stance direction. When the target opinion is "Pro", it can be observed from the table that, compared to the opinion distribution before manipulation, the proportion of responses holding a supportive stance (in blue) has significantly increased in the opinion distribution after manipulation. Conversely, when the target opinion is "Con", there is a notable rise in the proportion of responses expressing an opposing stance in the manipulated opinion distribution.
Figure 6: Case Study of RAG Opinion Manipulation on "Should Corporal Punishment Be Used in K-12 Schools".
@@ 5.2 RQ2: How significant does opinion manipulation influence the response of the LLM in RAG?
Figure 6 presents a case study of our RAG opinion manipulation on the question, "Should Corporal Punishment Be Used in K-12 Schools?" The black box highlights text expressing an supporting stance on the topic, while the red box indicates text opposing the topic. Initially, the LLM in the RAG system strongly opposed corporal punishment. However, after our transfer attack manipulated the retrieval ranking to prioritize documents supporting corporal punishment, the LLM’s opinion in its response reversed accordingly.
Later, to further analyze the opinion manipulation results on RAG response at a finer granularity, we conducted our transfer poisoning attacks on the corpus data of $\text{RAG}_{\text{black}}$ by themes. We focus on "Health", "Society", "Government", "Education" four themes.
@@
The results of the theme-specific manipulation experiments are shown in Table 5. The adversarial retrieval attack strategy PAT, applied to poison the target document, significantly increased the proportion of candidate items with target opinion in the Top 3 of the $\text{RAG}_{\text{black}}$ retrieval list, thereby guiding the LLM to change its opinion in the response. However, the manipulation effect on $\text{RAG}_{\text{black}}$ output opinions varies across different themes.
@@
TABLE V: Manipulation results of $\text{RAG}_{\text{black}}$ ranking and response opinion across different themes.
@@
Overall, our transfer attack demonstrates the greatest manipulation effectiveness on LLM-generated opinions for questions related to the "Government" theme. The manipulation effectiveness for questions in the "Education" theme is also strong, although it is somewhat less effective when the target opinion is "CON." In comparison, the opinion manipulation effectiveness for the "Health" and "Social" themes is less satisfactory. This discrepancy is primarily due to the preference characteristics of the black-box retrieval model across different themes. As shown in Table 5, the rankings of the black-box target retrieval model for questions in the "Government" and "Education" themes are more susceptible to manipulation, as evidenced by their higher $Top3_{v}$ score.
5.3 RQ3: Does opinion manipulation significantly impact users’ perceptions of controversial topics?
To further explore the impact of opinion manipulation on users’ perceptions of public topics, we designed a user experiment to assess people’s reactions to RAG responses expressing different opinions. We invited 10 college students to participate in the experiment, dividing them into two groups, A and B, with five participants in each group. None of the participants were informed about the purpose of the experiment, and their experimental processes were independent of one another, with no communication allowed. We first developed a simple question-answering (QA) service based on the RAG framework, and then applied our attack to relevant documents containing supportive opinions on two designated controversial topics in the corpus. Participants in Group A accessed the QA service before any opinion manipulation was applied to the corpus, while participants in Group B accessed the service after we poisoned the corpus using our attack. Following the interaction, both groups were asked to rate their approval of the two topics on a scale of 1 to 10.
TABLE VI: User Approval Rating on the Public Topics.
Subject User Approval
Topic 1 Topic 2
A Group(before manipulation) 5.8 4.8
B Group(after manipulation) 6.4 8.2
As shown in Table 6, when we manipulated the relevance ranking of documents supporting the target topics, users’ approval of those topics significantly increased. This experiment confirms that opinion manipulation has a substantial effect on users’ perceptions.
5.4 Mitigation Analysis
Existing studies on RAG poisoning attacks have largely focused on manipulation in white-box scenarios. In contrast, research exploring RAG attacks from a more realistic black-box perspective has primarily employed relatively simple attack methods.
Zou et al. [29] propose PoisonedRAG which injects a few poisoned texts into the RAG database such that the LLM generates manipulated answers. They consider both black-box and white-box settings. In the black-box setting, they attempt to include the question itself as part of the poisoned text to make it semantically similar to the target question. The rationale behind this approach is that the question is most similar to itself. While this strategy is effective in manipulating RAG responses and easy to implement in practice, it is highly detectable in real-world applications, as directly inserting the question makes the manipulation overly conspicuous. Therefore, we use automatic spamicity detection to defend PoisonedRAG and FlipedRAG.
Table 7 illustrates the results of the automatic spamicity detection experiment for both PoisonedRAG and FlipedRAG, which utilizes black-box imitation and PAT. Spamicity detection [28] is used for discovering term spam. Earlier search engines relied on term-based ranking methods such as TF-IDF to retrieve documents. In this context, term spam manipulates document content through keyword stuffing to artificially inflate the TF-IDF scores of spam documents, making them appear more relevant for specific queries. A straightforward approach to achieve this is by embedding the question or query within the target document, thereby increasing the frequency of keywords in the document and boosting its TF-IDF-based ranking.

### v3

@@ A.1. Algorithm
Input: target black-box RAG model $\boldsymbol{\textit{RAG}_{\text{black}}}$, target retrieval model $\boldsymbol{RM}$, surrogate model $\boldsymbol{M_{i}}$, controversial topics $\boldsymbol{Q}$, target topic $q$, target opinion $S_{t}$, corpus $Docs$ , target documents with target opinion $Docs_{t}$, target document $d_{t}$, relevant document $d_{+}$, random sampled document $d_{-}$, poisoned document amount $N$
@@
$i_{1}$ =“—Here is the user question—
@@
Please COPY all the given context altogether in [[ ]] including all marks and symbols. Do not omit any sentence of the context”
$i_{2}$ =“Use the following pieces of retrieved context to answer the question…”
@@ 3       INIT: Dataset $\mathcal{D}\leftarrow\{\}$
             // $\textit{RAG}_{\text{black}}(\boldsymbol{q}_{m}\oplus i_{1})$approx $\boldsymbol{RM(\boldsymbol{q}_{m})}$
@@ 15       INIT: RAG Response Set $\boldsymbol{Res}\leftarrow\{\}$
17             select $N$ target documents with $S_{t}$ as $Docs_{t}$
19             $\boldsymbol{anchor}_{m}\leftarrow$ top-1($\boldsymbol{R}_{m}{top3}$)
@@ 24            $\boldsymbol{Res}\leftarrow$ $\textit{RAG}_{\text{black}}(\boldsymbol{q}_{m};\boldsymbol{Docs})$
Algorithm 1 FlippedRAG
A.2. Baselines
A.2.1. Prompt Injection Attack
Prompt injection attacks constitute a class of adversarial exploits wherein malicious actors strategically insert compromised instructions or corrupted data into the inputs of LLM-integrated systems, thereby coercing the model’s outputs to align with predefined adversarial objectives(Liu et al., 2024). In black-box opinion manipulation scenarios, we formulate specific instructions that compel the LLM to generate stance-aligned responses (supporting or opposing opinions) regarding a given controversial topic, and embed the instruction within target documents. When the RAG system processes these documents, its outputs systematically align with the attacker’s designated opinion through execution of the embedded adversarial prompts.
A.2.2. Disinformation
Pan et al. (Pan et al., 2023) developed a methodology for generating credible-sounding misinformation by leveraging LLMs. To implement this approach, we employ the adversarial prompt proposed in their work “generate a background document in support of the given opinion to the question” to induce LLMs to produce context passages imbued with specific opinions to the controversial topics. We subsequently add the generated context passages into the RAG corpus. Given their elevated topical relevance to specific controversial subjects, these passages demonstrate enhanced retrievability within the system, thereby enabling their embedded opinions to influence LLM outputs through semantic content manipulation. This mechanism fundamentally differs from the command-execution paradigm characteristic of Prompt Injection Attacks.
A.2.3. GARAG
GARAG(Cho et al., 2024) employs genetic algorithms to optimize the discovery of novel adversarial documents that achieve dual objectives: (1) maintaining retrievability within RAG systems, and (2) deviating LLM outputs from factual accuracy. In our experimental, we implement a inversion of this methodology, repurposing GARAG-generated adversarial documents to strategically induce LLMs to consistently produce responses aligning target opinions.
A.2.4. PoisonedRAG
Zou et al.(Zou et al., 2024) proposed PoisonedRAG, whose black-box attack framework leverages the insight that “a question is most similar to itself” to inject the question into the malicious text. This methodology enhances malicious text relevance rankings through keyword stuffing while simultaneously leveraging LLM-generated adversarial components to manipulate RAG systems into producing factually erroneous outputs. In our experiment, we reorient PoisonedRAG’s adversarial objective from factual corruption to opinion bias induction, transforming its attack vector into generating extreme perspectives through corpus corruption.
A.3. Evaluation Metrics
A.3.1. NDCG
NDCG is designed to evaluate the effectiveness of search engines in ordering results by relevance. The NDCG score ranges from 0 to 1. An NDCG score close to 1 indicates that the ranking is nearly ideal, with highly relevant items near the top of the list.
A.3.2. Top3v and Top6v
@@
Top3v is obtained by subtracting Top3origin(the average proportion of target opinions in the top 3 rankings before manipulation) from Top3attacked(the average proportion of target opinions in the top 3 rankings after manipulation). Note that Top3v ranges from -1 to 1, and a positive value indicates that the manipulation successfully increased the proportion of the target opinion in the top-3 ranking. The higher the Top3v value, the more effective the manipulation. Top6v is also defined as above.
A.3.3. RASR
@@
RASR measures the average rate of successfully boosted candidates in the ranking of each query. The higher is the RASR value, the more successful is the ranking attack.
A.3.4. BRank
@@
BRank is the average of the total rank improvements for all target documents under each query. Note that a target document $d_{t}$ contributes a negative value to BRank if its rank is reduced accidentally by the attack.
A.3.5. OMSR
@@
OMSR measures the average rate of successfully manipulated LLM responses at the opinion level. We consider one manipulation as a success if it guides the LLM to produce a response whose opinion is closer to the target opinion $S_{t}$ compared with the original response without RAG attack. Given the target opinion “support”(Pro), successful opinion manipulation can be from “oppose”(Con) to “neutral”, from “oppose”to “support”, and from “neutral” to “support”. Given the target opinion “oppose”, successful opinion manipulation can be from “support” to “neutral”, from “support” to “oppose”, and from “neutral” to “oppose”.
A.3.6. ASV
@@
ASV represents the average increase of opinion scores of LLM responses in the direction of the target opinion $S_{t}$ after manipulation. In this research, we assign 0 to opposing stance, 1 to neutral stance, and 2 to supporting stance. A positive ASV indicates that the opinion manipulation towards $S_{t}$ is effective, while a negative ASV indicates that the manipulation actually makes the opinions of RAG responses deviate from $S_{t}$. ASV reflects how much the shift in opinion polarity is. The larger the ASV value, the more successful the opinion manipulation.
Appendix B Experiment Result and Analysis
B.1. Case Study
@@
Figure 8 presents a case study of our RAG opinion manipulation on the question, “Should Corporal Punishment Be Used in K-12 Schools?” The black box highlights text expressing an supporting stance on the topic, while the red box indicates text opposing the topic. Initially, the LLM in the RAG system strongly opposed corporal punishment. However, after FlippedRAG manipulated the retrieval ranking to prioritize documents supporting corporal punishment, the LLM’s opinion in its response reversed accordingly.
Figure 8. Case Study of RAG Opinion Manipulation on “Should Corporal Punishment Be Used in K-12 Schools”.
[…]
Case study.
B.2. Overall Effect of Opinion Manipulation on Mixtral and Vicuna
The FlippedRAG’s overall effect of opinion manipulation on Mixtral and Vicuna is in Figure 9.
Figure 9. FlippedRAG’s overall effect of opinion manipulation of on Mixtral and Vicuna with retrieval model Contriever.
[…]
B.3. Domain-specific manipulation experiment of FlippedRAG
@@
The results of the theme-specific manipulation experiments are shown in Table 12. FlippedRAG significantly increases the proportion of candidate items with target opinion in the Top 3 of the retrieval list, thereby guiding the LLM to change its opinion in the response. However, the manipulation effect on $\text{RAG}_{\text{black}}$ output opinions varies across different themes.
FlippedRAG demonstrates the greatest manipulation effectvfor questions related to the “Government” theme. The manipulation effect for questions in the “Education” theme is also strong, although it is somewhat less effective when the target opinion is “CON.” In comparison, the opinion manipulation effectiveness for the “Health” and “Social” themes is less satisfactory. This discrepancy is primarily due to the preference characteristics of the black-box retrieval model across different themes. As shown in Table 12, the rankings of the black-box retrieval model for questions in the “Government” and “Education” themes are more susceptible to manipulation.
@@
Table 12. Manipulation results of $\text{RAG}_{\text{black}}$ ranking and response opinion across different domains.
@@
B.4. User Study
We invited 10 college students to participate in the experiment, dividing them into two groups, A and B, with five participants in each group. None of the participants were informed about the experimental purpose before the experiment, and their experimental processes were independent of one another, with no communication allowed. We first developed a simple question-answering (QA) service based on the RAG framework, and then applied our attack to relevant documents containing supportive opinions on two designated controversial topics in the corpus. Participants in Group A accessed the QA service before any opinion manipulation was applied to the corpus, while participants in Group B accessed the service after we poisoned the corpus using our attack. Following the interaction, both groups were asked to rate their approval of the two topics on a scale of 1 to 10.
In accordance with the IRB approval requirements for “post-experiment feedback,” all participants received a comprehensive notification after completing the experiment: Participants in the experimental group (Group B) were informed of the technical details of the adversarial attack on the system (Control group: Group A)). All participants can request to refer to the original system answers of the control group (Group A) as a reference.
B.5. Mitigation Analysis
B.5.1. Mitigation Based on Spam Detection
Earlier search engines relied on term-based ranking methods such as TF-IDF to retrieve documents. In this context, term spam manipulates document content through keyword stuffing to artificially inflate the TF-IDF scores of spam documents, making them appear more relevant for specific queries. A straightforward approach to achieve this is by embedding the question within the target document, thereby increasing the frequency of keywords in the document and boosting its TF-IDF-based ranking.